# Limits on the Axion–Photon Coupling from Chandrayaan-2

### v1

@@ Abstract
Axions and axion-like particles (ALPs) have gained immense attention in searches for beyond Standard Model (BSM) physics. Experiments searching for axions leverage their predicted couplings to Standard Model (SM) particles to look for observable signals. Though weak, these couplings allow axions to be produced abundantly in the interiors of stars such as the Sun. Once created, axions can escape the Sun and while passing through the solar atmosphere, oscillate into photons in the magnetic field producing x-rays. For the first time, we used data from the observation of soft x-rays from the quiet Sun during the 2019-20 solar minimum by the solar x-ray monitor (XSM), onboard India’s Chandrayaan-2 lunar exploration mission, to constrain the coupling of axions to photons ($g_{a\gamma\gamma}$). Using the latest models of the solar atmosphere to calculate the magnetic field and plasma frequency, we constrain $g_{a\gamma\gamma}\lesssim(0.47\,-\,2.2)\times 10^{-10}$ GeV-1 at 95% confidence level for axion masses $m_{a}\lesssim 5\times 10^{-4}$eV.
I INTRODUCTION
@@
Axions and axion-like particles (ALPs) are among the most compelling candidates for physics beyond the Standard Model (SM) of particle physics. Such particles naturally arise in several BSM theories as a pseudo-Nambu-Goldstone boson (pNGB) of some spontaneously broken approximate global $U(1)$ symmetry [1, 2, 3]. The most famous among such particles is the QCD axion, which arises in the Peccei-Quinn (PQ) model [4, 5, 6, 7, 8], due to the breaking of the $U(1)_{\rm PQ}$ symmetry. In this article, we use the term axions for generic pNGBs associated with nonlinearly realized approximate global symmetries.
Due to a plethora of theories suggesting their existence, a large number of experiments have been proposed or are underway to search for these axions over a wide range of masses stretching from $10^{-22}$ eV upto several GeV. The theories that predict the existence of axions also suggest the existence of extremely weak but non-negligible interactions between axions and the SM particles. Due to these weak interactions, light axions ($m_{a}\lesssim$ few eV) are also excellent candidates for cold dark matter (DM). Most of the present and future experiments such as ABRACADABRA [9, 10], DMRadio [11, 12, 13], ADMX [14, 15, 16], HAYSTAC [17] and MADMAX [18, 19] looking for light axions are designed to search for these low mass axion DM by utilizing their predicted coupling to photons. However, the mere existence of axions in the BSM spectrum does not imply that they have to be DM. In that case, we need to consider other strategies to search for axions.
Due to their couplings with the SM particles, axions can also be produced in a variety of different astrophysical systems. For example, the large temperatures, densities, and volumes of stars and other stellar objects make them efficient factories of such weakly-coupled axions. For sufficiently small couplings, these axions will free-stream out to infinity. These axions can be detected using different techniques. A prime target for such search strategies is our Sun, a main-sequence star as well as the closest star to Earth. In the core of the Sun, because of the high temperature ($T\sim\rm few\>\;keV$) and high density ($\rho\sim 150\,\text{gm}\,\text{cm}^{-3}$), axions can be produced copiously. These solar axions, having energies in the range of a few keV, escape the sun and travel towards Earth. Hence, the Earth is expected to be bathed with a large flux of solar axions. The CERN Axion Solar Telescope (CAST) has been searching for these solar axions. In the absence of any positive signal, the CAST experiment has placed stringent bounds in the axion parameter space [20, 21, 22]. Furthermore, the International AXion Observatory (IAXO) has been proposed as a successor to the CAST experiment to search for these solar axions with an enhanced sensitivity [23].
@@
In this work, we explore a novel avenue for detecting solar axions. During the 2019-2020 solar minimum, when solar activity was at its lowest, the Chandrayaan-2 lunar orbiter of the Indian Space Research Organization (ISRO) measured the soft x-ray spectrum originating from the Sun using the solar x-ray Monitor (XSM) on board [24, 25]. This soft x-ray spectrum consists of two components - the background quiet-Sun emission and the emission from the flaring plasma. Now if axions exist and are produced inside the Sun, these solar axions (with energies in the range of few keV) can also contribute to the total solar soft x-ray emission due to the conversion of axions into photons in the solar magnetic field. In this work we conduct a search for axion-induced x-ray emission from the Sun using the observed soft x-ray emission data of the quiet Sun and we put constraints on the axion-photon coupling. We find no evidence of an axion-induced signal over the background and consequently set an upper bound on the axion-photon coupling, which is comparable to the present constraints from the CAST experiment and is limited only by the collecting area of the XSM instrument.
@@
A similar analysis using data from the NuSTAR observation of the Sun during the solar minimum has been presented in [26]. However, unlike NuSTAR, which was able to observe only a portion of the Sun because of its small field of view (FOV) ($12^{\prime}\times 12^{\prime}$), the XSM observes emission from the entire solar disk. Our work provides a compelling case for having a dedicated solar observatory operational during the next solar minimum which can significantly improve upon these bounds in an independent fashion.
The paper is structured as follows. In Section II we outline the production mechanism of axions inside the Sun. In Section III we study the axion-photon conversion in the solar magnetic field and calculate the x-ray signal originating from it. In Section IV, we describe the data that we have used for our analysis and in Section V we present the data analysis procedure that we have adopted in our work. Finally, we present our results in Section VI.
II AXION PRODUCTION INSIDE THE SUN
@@
We start with the following Lagrangian for the axion field $a$:
@@
Following [27] we calculate the axion emission rate per unit volume and we integrate it over a standard solar model [28] and finally dividing by the total surface area of the Sun, we obtain the flux of axions coming out of the Sun. For the light axions ($m_{a}<<\rm eV$) that we shall consider in this work, the spectral axion fluxes (in units of cm-2 s-1 keV-1) produced by the Sun via the four processes are well approximated by the following fitting formulae [21]
@@
Fig. 1 shows the fluxes of solar axions at the solar surface produced via the Primakoff process (solid red line) and through the combined Bremsstrahlung and Compton processes (solid blue line), for $g_{a\gamma\gamma}=10^{-10},\text{GeV}^{-1}$ and $g_{aee}=1.3\times 10^{-13}$. This value of $g_{aee}$ is adopted throughout the analysis. The total solar axion flux is obtained as the sum of the individual contributions from the Primakoff, Bremsstrahlung, and Compton processes, and therefore depends on both $g_{a\gamma\gamma}$ and $g_{aee}$. Within the parameter space and energy range considered here, the Primakoff-induced flux is substantially larger than the combined Bremsstrahlung and Compton components. Nevertheless, for completeness, all production channels are included in the subsequent analysis.
III AXION-PHOTON CONVERSION OUTSIDE THE SUN
Relativistic axions produced in the solar core propagate along nearly radial trajectories and emerge from the solar surface. As they traverse the solar magnetic field, a fraction convert into photons, producing x-rays. The differential solar x-ray flux resulting from axion–photon conversion in the solar atmosphere is given by
@@ III AXION-PHOTON CONVERSION OUTSIDE THE SUN
To calculate the conversion probability, we start with the axion-photon mixing equation in an external magnetic field [29]
@@
$M(\ell)=\begin{pmatrix}-\frac{\omega_{p}^{2}(\ell)}{2\omega}&0&0\\ 0&-\frac{\omega_{p}^{2}(\ell)}{2\omega}&g_{a\gamma\gamma}\frac{B_{T}(\ell)}{2}\\ 0&g_{a\gamma\gamma}\frac{B_{T}(\ell)}{2}&-\frac{m_{a}^{2}}{2\omega}\\ \end{pmatrix}\,$. (5)
Here $\omega_{p}(\ell)=\sqrt{4\pi\,\alpha\,n_{e}(\ell)/m_{e}}$ is the plasma frequency in the solar atmosphere, where $\alpha$ is the fine structure constant. For this work, we neglect the Euler-Heisenberg term and Faraday rotation related to the longitudinal component of the external field. Since the axion field only couples to $A_{\parallel}$, axion-photon conversion can be described by the bottom $2\times 2$ sub-block of $M$, which we denote as $\hat{M}$. Starting from the point $\ell=0$, the solution of Eq. (4) at some location $\ell=h$ is given by the path-ordered transfer matrix [29, 30]
@@
The above integral can be evaluated perturbatively in powers of $g_{a\gamma\gamma}$ [29]. Working up to first order in perturbation theory, one can find that the probability of conversion of axions into photons after traveling a distance $h$ from the solar surface in the solar magnetic field is given by,
$P_{a\gamma}(h)=$ $\,|\langle A_{\parallel}(h)|a(0)\rangle|^{2}$ (7)
$=$ $\,\dfrac{1}{4}\,g_{a\gamma\gamma}^{2}\,e^{-\int_{0}^{h}\,d\ell\,\Gamma(\ell)}\,\left|\int_{0}^{h}\,d\ell\,B_{T}(\ell)\,\right$.
$\left.\times\,e^{i\int_{0}^{\ell}d\ell^{\prime}q(\ell^{\prime})}\,e^{\frac{1}{2}\int_{0}^{\ell}d\ell^{\prime}\Gamma(\ell^{\prime})}\right|^{2},$
where $q(\ell)=(\omega_{p}^{2}(\ell)-m_{a}^{2})/2E_{a}$ is the momentum exchanged between the photon and the axion, both carrying energy $E_{a}$ and $\Gamma(\ell)$ is the coefficient of absorption of x-ray in the solar atmosphere.
@@
In this work, we compute the conversion probability numerically by integrating the right hand side of Eq. (7) up to the required height $h$ above the surface of the Sun. To do so, we utilize the magnetic field profile and plasma frequency profile of the quiet Sun used in Ref. [26]. The absorption coefficient on the other hand is computed as $\Gamma=\Sigma_{Z}\,n_{Z}\,\sigma_{Z}$ where $n_{Z}$ is the number density of element of atomic number $Z$ and $\sigma_{Z}$ is the total attenuation cross-section of a photon passing through a gas of atomic number $Z$. To compute $\Gamma$, we consider all elements up to atomic number $Z=30$. The individual elemental abundance in the solar atmosphere is obtained from the CHIANTI atomic database [31, 32] by using the python interface ChiantiPy [33]. For each element, we take the total photon attenuation cross-section with coherent scattering from the XCOM database by NIST [34].
The top panel of Fig. 2 illustrates the variation of the axion–photon conversion probability, $P_{a\gamma}$, in the solar atmosphere as a function of the height above the solar surface, $h$, for a range of axion masses $m_{a}\in\{10^{-3},10^{-4},10^{-5},10^{-6},10^{-7}\}\,\mathrm{eV}$. The calculation is performed for an axion–photon coupling $g_{a\gamma\gamma}=10^{-10}\,\mathrm{GeV^{-1}}$ and an axion energy $E_{a}=4\,\mathrm{keV}$. As illustrated in the figure, the conversion probability $P_{a\gamma}$ asymptotically approaches a constant value at large distances from the Sun, with its magnitude increasing as the axion mass decreases from $10^{-3}\,\mathrm{eV}$ to $10^{-7}\,\mathrm{eV}$. For even smaller masses ($m_{a}\lesssim 10^{-7}\,\mathrm{eV}$), the probability becomes nearly independent of $m_{a}$, leading to overlapping curves. Therefore, in this low-mass regime, the resulting constraints on $g_{a\gamma\gamma}$ are expected to be largely insensitive to the axion mass, as discussed later.
The bottom panel of Fig. 2 shows the variation of the axion–photon conversion probability, $P_{a\gamma}$, evaluated far from the Sun, as a function of the axion mass $m_{a}$ and axion energy $E_{a}$, for a fixed coupling $g_{a\gamma\gamma}=10^{-10}\,\mathrm{GeV^{-1}}$. It is evident that the conversion probability remains nearly independent of the axion energy within the range of energies considered. Consequently from Eq. (3), the resulting x-ray signal produced in the solar atmosphere through axion conversion is expected to exhibit the same spectral shape as the solar axion flux shown in Fig. 1, differing only by an overall normalization factor determined by the conversion probability.
IV CHANDRAYAAN-2 OBSERVATION OF THE QUIET SUN
@@
Solar X-ray Monitor (XSM) on-board Chandrayaan-2 measures the disk-integrated solar x-ray spectrum in 1 – 15 keV energy range [35, 36]. Here we use the XSM observations during its first two observing seasons from September 2019 to May 2020, when the solar activity was at its minimum. Data taken during the same observing periods of the quiet Sun when no active regions or microflares are present, as reported in [24], are used for the analysis to obtain the total quiet Sun x-ray spectrum.
@@
In addition to emission from the Sun, XSM also records background which primarily consists of Cosmic X-ray Background (CXB), charged particle background, and emission from other bright astrophysical sources. This background component needs to be subtracted from the observations to obtain residual x-ray emission from the Sun. While the CXB component is steady, the measured CXB by XSM varies. Some part of the XSM FOV is usually obstructed by the Moon blocking the CXB emission, and the blocked fraction varies due to the variation in the orbital altitude of the spacecraft. The charged particle background has temporal variations due to passage through geotail regions, solar energetic particle (SEP) events, Coronal Mass Ejection (CME) events, which are also recorded by the upper level discriminator (ULD) counts of the XSM instrument [36]. As the spacecraft attitude varies, bright astrophysical sources, Sco X-1 being the brightest source in this energy range, enter within the XSM FOV for some periods and contribute to the background. For the purpose of the present analysis, such periods of varying background with high ULD counts and presence of the Sco X-1 within the FOV are ignored while generating the quiet Sun x-ray spectrum to ensure that the background is as clean as possible. We use the modules of XSM Data Analysis Software [37] with custom good time intervals to generate the quiet Sun spectrum, resulting in the net effective exposure of $3.72\times\mathrm{10^{6}}$ s.
@@
Fig. 3(a) shows the XSM quiet Sun spectrum in the 1–15 keV energy range, thus obtained along with a background spectrum when the Sun is out of FOV. The XSM spectrum of the quiet Sun is dominated by the quiescent coronal emission in the low energies up to $\sim$3 keV[24] and thus we use the spectrum above 3.4 keV to explore the x-ray emission produced by axion conversion. To obtain the most conservative residual x-ray emission from the Sun, the CXB component alone is subtracted from this observed XSM spectrum. CXB model spectrum [38] taking into account the fraction of FOV obstructed by the Moon is subtracted from the observed XSM spectrum, and the resultant spectrum is shown in Fig. 3(b) top panel. The uncertainties shown are 1-sigma and arise from Poisson errors on the observed counts. The slight excess near 7.5 keV is due to a known instrumental line and may be ignored [36]. As this spectrum includes the contribution from the charged particle background and other sources, a more realistic residual solar x-ray spectrum can be obtained by subtracting the background observed when the Sun is out of the FOV from the quiet Sun observations. However, this requires ignoring the periods when other bright sources contribute to the background. Sources that were bright enough to have a signal in XSM during the observing period were identified from the MAXI all-sky monitor data, and periods when such 13 sources were within the XSM FOV are ignored, resulting in an effective quiet Sun exposure of $2.15\times\mathrm{10^{6}}$ s. The resulting spectrum after subtracting the background is shown in the middle panel of Fig. 3(b) (Case 2), errors incorporate uncertainties on the observed quiet Sun spectrum and the background spectrum. We also consider Case 3, where the entire observed spectrum in Case 1 is assumed to arise from the background, and if the background were perfectly modeled, the resultant spectrum obtained is shown in the bottom panel of Fig. 3(b). This is obtained by empirically modeling the observed spectrum and subtracting the best-fit model from the observed spectrum.
V DATA ANALYSIS
@@
The expected number of photons per unit time per unit energy in a given energy bin $i$ of the XSM solely due to axion-photon conversion in the solar atmosphere is
$N^{a}_{\gamma,\,\rm exp}(E_{i})=\dfrac{R_{\odot}^{2}}{(\text{1 A.U.})^{2}}\,A_{\rm eff}(E_{i})\,\dfrac{d\phi_{a}}{dE_{a}}(E_{i})\,P_{a\gamma}(E_{i}),$ (8)
@@
where $E_{i}$ is the energy of the $i$-th energy bin, $A_{\rm eff}(E_{i})$ is the effective collecting area of the XSM in the $i$-th energy bin which we assume to be constant and equal to $0.32\;\text{mm}^{2}$ for all bins and $R_{\odot}$ is the radius of the Sun. Given the observed number of counts per unit time per unit energy by the XSM, $N_{\gamma,\,obs}(E_{i})$, for a particular background subtraction scheme, we compute the likelihood assuming independent Gaussian statistic in each energy bin:
@@
$\mathcal{L}(g_{a\gamma\gamma};m_{a})\propto\prod_{i}\,\exp\left[-\dfrac{1}{2}\left(\dfrac{N^{a}_{\gamma,\,\rm exp}(E_{i})\,-\,N_{\gamma,\,obs}(E_{i})}{\sigma(E_{i})}\right)^{2}\right],$ (9)
@@
where $\sigma(E_{i})$ denotes the uncertainty in the observed counts in the $i$-th energy bin. To compute our limits we consider all the energy bins between $3.4$ keV and $15$ keV.
@@
To derive constraints, we fix the axion mass $m_{a}$ and adopt a uniform prior on the axion–photon coupling $g_{a\gamma\gamma}$ over the range $10^{-14},10^{-8}]\ \mathrm{GeV}^{-1}$. Following a Bayesian approach, the posterior probability density function for $g_{a\gamma\gamma}$, denoted as $\mathcal{P}(g_{a\gamma\gamma}\,|\,\text{data})$, is obtained using the PyMultiNest package, which uses the MultiNest algorithm [39] to efficiently sample the likelihood and normalize the posterior. The 95% credible upper limit on $g_{a\gamma\gamma}$ is then determined by finding the value of $g_{a\gamma\gamma}$ below which 95% of the posterior probability is contained. This procedure is repeated for axion masses in the range $10^{-12},10^{-3}]\ \mathrm{eV}$, yielding the corresponding 95% credible upper limits on $g_{a\gamma\gamma}$ as a function of $m_{a}$.

### v2

@@ Abstract
Axions and axion-like particles (ALPs) have gained immense attention in searches for beyond Standard Model (BSM) physics. Experiments searching for axions leverage their predicted couplings to Standard Model (SM) particles to look for observable signals. Though weak, these couplings allow axions to be produced abundantly in the interiors of stars such as the Sun. Once created, axions can escape the Sun and while passing through the solar atmosphere, oscillate into photons in the magnetic field producing x-rays. For the first time, we used data from the observation of soft x-rays from the quiet Sun during the 2019-20 solar minimum by the solar x-ray monitor (XSM), onboard India’s Chandrayaan-2 lunar exploration mission, to constrain the coupling of axions to photons ($g_{a\gamma\gamma}$). Using the latest models of the solar atmosphere to calculate the magnetic field and plasma frequency, we constrain $g_{a\gamma\gamma}\lesssim(0.50\,-\,2.26)\times 10^{-10}$ GeV-1 at 95% confidence level for axion masses $m_{a}\lesssim 5\times 10^{-4}$eV.
@@
Introduction. Axions and axion-like particles (ALPs) are well-motivated candidates for physics beyond the Standard Model (SM). They arise generically as pseudo-Nambu–Goldstone bosons (pNGBs) of spontaneously broken approximate global $U(1)$ symmetries in many extensions of the SM [Kim:2008hd, DiLuzio:2020wdo, Arvanitaki:2009fg]. The best-known example is the QCD axion, originating from the Peccei–Quinn (PQ) solution to the strong CP problem [Peccei:1977hh, Peccei:1977ur, Weinberg:1977ma, Wilczek:1977pj, Peccei:2006as]. Throughout this work, we use the term axions to denote generic pNGBs associated with nonlinearly realized approximate global symmetries.
Axions are predicted to have masses ranging, from $10^{-22}\,\mathrm{eV}$ to several GeV, and interact extremely weakly with SM particles. Consequently, light axions ($m_{a}\lesssim\mathrm{few}\,\mathrm{eV}$) are compelling cold dark matter (DM) candidates [Preskill:1982cy, Abbott:1982af, Dine:1982ah, Khlopov:1999tm]. Hence, most experimental searches for light axions target axion DM through its coupling to photons [Ouellet:2018beu, Ouellet:2019tlz, Silva-Feaver:2016qhh, DMRadio:2022jfv, DMRadio:2022pkf, ADMX:2003rdr, ADMX:2018gho, ADMX:2019uok, HAYSTAC:2018rwy, Caldwell:2016dcw, Millar:2016cjp]. However, axions need not constitute the DM, motivating complementary search strategies.
Astrophysical objects such as stars provide powerful laboratories to search for axions. Owing to their high temperatures, densities, and volumes, they can efficiently produce axions [Vysotsky:1978dc, Raffelt:2006cw, Lella:2022uwi, Lella:2023bfb, Carenza:2024ehj, Lecce:2025dbz, Fiorillo:2025gnd], which can freely stream out of the stars for sufficiently small couplings. In particular, the Sun is a prime target: axions produced in its core at $T\sim\mathrm{few}\,\mathrm{keV}$ and $\rho\sim 150\,\mathrm{g\,cm^{-3}}$ escape with $\mathcal{O}($keV)) energies. The CERN Axion Solar Telescope (CAST) has searched for such solar axions and, in the absence of a signal, has placed stringent bounds on axion parameter space [CAST:2004gzq, Barth:2013sma, CAST:2017uph].
@@
In this work, we explore a novel avenue for detecting solar axions by using the solar soft x-ray spectrum measured by the solar x-ray Monitor (XSM) on board the Chandrayaan-2 lunar orbiter of the Indian Space Research Organization (ISRO), during the 2019-2020 solar minimum [2021ApJ...912L..12V, Vadawale:2021pis]. This soft x-ray spectrum consists of two components - the background quiet-Sun emission and the emission from the flaring plasma. Additionally, solar axions will also contribute to the total solar soft x-ray emission due to their conversion into photons in the solar magnetic field. We conduct a search for this axion-induced x-ray emission from the Sun using the observed solar soft x-ray spectrum and put constraints on the axion-photon coupling. We find no evidence of axion-induced signal over the background and consequently set an upper bound on the axion-photon coupling, which is comparable to the present constraints from the CAST experiment and is limited only by the collecting area of the XSM.
@@
A similar analysis using data from the NuSTAR observation of the Sun during the solar minimum has been presented in [Ruz:2024gkl]. However, unlike NuSTAR, which was able to observe only a portion of the Sun because of its small field of view (FOV) ($12^{\prime}\times 12^{\prime}$), the XSM observes emission from the entire solar disk. Our work provides a compelling case for having a dedicated solar observatory operational during the next solar minimum which can significantly improve upon these bounds in an independent fashion.
@@
Axion production inside the Sun. We start with the following Lagrangian for the axion field $a$:
@@
Following [Raffelt:1985nk] we calculate the axion emission rate per unit volume and we integrate it over a standard solar model [Bahcall:2004pz] and finally dividing by the total surface area of the Sun, we obtain the flux of axions coming out of the Sun. For the light axions ($m_{a}<<\rm eV$) that we shall consider in this work, the spectral axion fluxes (in units of cm-2 s-1 keV-1) produced by the Sun via the four processes are well approximated by the following fitting formulae [Barth:2013sma]
@@
Fig. 1 shows the fluxes of solar axions at the solar surface produced via the Primakoff process (solid red line) and through the combined Bremsstrahlung and Compton processes (solid blue line), for $g_{a\gamma\gamma}=10^{-10}\,\text{GeV}^{-1}$ and $g_{aee}=1.3\times 10^{-13}$. The total solar axion flux is obtained as the sum of the individual contributions from the Primakoff, Bremsstrahlung, and Compton processes, and therefore depends on both $g_{a\gamma\gamma}$ and $g_{aee}$. Within the parameter space and energy range considered here, the Primakoff-induced flux is substantially larger than the combined Bremsstrahlung and Compton components. Nevertheless, for completeness, all production channels are included in the subsequent analysis.
Axion-photon conversion outside the Sun. Relativistic axions produced in the solar core propagate along nearly radial trajectories and emerge from the solar surface. As they traverse the solar magnetic field, a fraction convert into photons, producing x-rays. The differential solar x-ray flux resulting from axion–photon conversion in the solar atmosphere is given by
@@
To calculate the conversion probability, we start with the axion-photon mixing equation in an external magnetic field [Raffelt:1987im]
@@
$M(\ell)=\begin{pmatrix}-\frac{\omega_{p}^{2}(\ell)}{2\omega}-i\dfrac{\Gamma(\ell)}{2}&0&0\\ 0&-\frac{\omega_{p}^{2}(\ell)}{2\omega}-i\dfrac{\Gamma(\ell)}{2}&g_{a\gamma\gamma}\frac{B_{T}(\ell)}{2}\\ 0&g_{a\gamma\gamma}\frac{B_{T}(\ell)}{2}&-\frac{m_{a}^{2}}{2\omega}\\ \end{pmatrix}\,$. (5)
Here $\omega_{p}(\ell)=\sqrt{4\pi\,\alpha\,n_{e}(\ell)/m_{e}}$ is the plasma frequency in the solar atmosphere, with $\alpha$ being the fine structure constant and $\Gamma(\ell)$ is the coefficient of absorption of x-ray in the solar atmosphere [vanBibber:1988ge]. For this work, we neglect the Euler-Heisenberg term and Faraday rotation related to the longitudinal component of the external field. Since the axion field only couples to $A_{\parallel}$, axion-photon conversion can be described by the bottom $2\times 2$ sub-block of $M$, which we denote as $\hat{M}$. Starting from the point $\ell=0$, the solution of Eq. (4) at some location $\ell=h$ is given by the path-ordered transfer matrix [Raffelt:1987im, Manzari:2024jns]
@@
The above integral can be evaluated perturbatively in powers of $g_{a\gamma\gamma}$ [Raffelt:1987im]. Working up to first order in perturbation theory, one can find that the probability of conversion of axions into photons after traveling a distance $h$ from the solar surface in the solar magnetic field as,
$P_{a\gamma}(h)=$ $\,|\langle A_{\parallel}(h)|a(0)\rangle|^{2}=\dfrac{1}{4}\,g_{a\gamma\gamma}^{2}\,e^{-\int_{0}^{h}\,d\ell\,\Gamma(\ell)}\,$
$\times\left|\int_{0}^{h}\,d\ell\,B_{T}(\ell)\,e^{i\int_{0}^{\ell}d\ell^{\prime}q(\ell^{\prime})}\,e^{\frac{1}{2}\int_{0}^{\ell}d\ell^{\prime}\Gamma(\ell^{\prime})}\right|^{2},$ (7)
where $q(\ell)=(\omega_{p}^{2}(\ell)-m_{a}^{2})/2E_{a}$ is the momentum exchanged between the photon and the axion, both carrying energy $E_{a}$.
@@
In this work, we compute the axion–photon conversion probability by numerically integrating Eq. (Limits on the axion–photon coupling from Chandrayaan-2 observations) up to the required height $h$ above the solar surface, using the magnetic-field and plasma-frequency profiles of the quiet Sun from Ref. [Ruz:2024gkl]. The absorption coefficient is computed as $\Gamma=\sum_{Z=1}^{30}n_{Z}\,\sigma_{Z}$, where $n_{Z}$ is the number density of element of atomic number $Z$ and $\sigma_{Z}$ is the total attenuation cross-section of a photon passing through a gas of atomic number $Z$. To compute $\Gamma$, we consider all elements up to atomic number $Z=30$. The individual elemental abundance in the solar atmosphere is obtained from the CHIANTI atomic database [2024ApJ...974...71D, 2021ApJ...909...38D] by using the python interface ChiantiPy [ChiantiPy:2022xx]. For each element, we take the total photon attenuation cross-section with coherent scattering from the XCOM database by NIST [xcom].
The top panel of Fig. 2 shows the axion–photon conversion probability $P_{a\gamma}$ as a function of height above the solar surface $h$ for various axion masses. The probability saturates at large distances and increases for lighter axions, becoming mass-independent for $m_{a}\lesssim 10^{-7}\,\mathrm{eV}$, where the curves converge.
The bottom panel of Fig. 2 shows $P_{a\gamma}$ far from the Sun as a function of $m_{a}$ and $E_{a}$. The probability is almost independent of axion energy within the relevant range. Hence, from Eq. (3), the resulting x-ray spectrum from axion conversion mirrors the solar axion spectrum in Fig. 1, differing only by an overall normalization set by the conversion probability.
@@
Chandrayaan-2 observation of the quiet Sun. Solar X-ray Monitor (XSM) on-board Chandrayaan-2 measures the disk-integrated solar x-ray spectrum in 1 – 15 keV energy range [2020CSci..118...45S, 2020SoPh..295..139M]. Here we use the XSM observations during its first two observing seasons from September 2019 to May 2020, when the solar activity was at its minimum. Data taken during the same observing periods of the quiet Sun when no active regions or microflares are present, as reported in [2021ApJ...912L..12V], are used for the analysis to obtain the total quiet Sun x-ray spectrum.
@@
In addition to emission from the Sun, XSM also records background which primarily consists of Cosmic X-ray Background (CXB), charged particle background, and emission from other bright astrophysical sources. This background component needs to be subtracted from the observations to obtain residual x-ray emission from the Sun. While the CXB component is steady, the measured CXB by XSM varies. Some part of the XSM FOV is usually obstructed by the Moon blocking the CXB emission, and the blocked fraction varies due to the variation in the orbital altitude of the spacecraft. The charged particle background has temporal variations due to passage through geotail regions, solar energetic particle (SEP) events, Coronal Mass Ejection (CME) events, which are also recorded by the upper level discriminator (ULD) counts of the XSM instrument [2020SoPh..295..139M]. As the spacecraft attitude varies, bright astrophysical sources, Sco X-1 being the brightest source in this energy range, enter within the XSM FOV for some periods and contribute to the background. For the purpose of the present analysis, such periods of varying background with high ULD counts and presence of the Sco X-1 within the FOV are ignored while generating the quiet Sun x-ray spectrum to ensure that the background is as clean as possible. We use the modules of XSM Data Analysis Software [2021A&C....3400449M] with custom good time intervals to generate the quiet Sun spectrum, resulting in the net effective exposure of $3.72\times\mathrm{10^{6}}$ s.
@@
Fig. 3(a) shows the XSM quiet Sun spectrum in the 1–15 keV energy range, thus obtained along with a background spectrum when the Sun is out of FOV. The XSM spectrum of the quiet Sun is dominated by the quiescent coronal emission in the low energies up to $\sim$3 keV[2021ApJ...912L..12V] and thus we use the spectrum above 3.4 keV to explore the x-ray emission produced by axion conversion. To obtain the most conservative residual x-ray emission from the Sun, the CXB component alone is subtracted from this observed XSM spectrum. CXB model spectrum [2010A&A...512A..49T] taking into account the fraction of FOV obstructed by the Moon is subtracted from the observed XSM spectrum, and the resultant spectrum is shown in Fig. 3(b) top panel. The uncertainties shown are 1-sigma and arise from Poisson errors on the observed counts. The slight excess near 7.5 keV is due to a known instrumental line and may be ignored [2020SoPh..295..139M]. As this spectrum includes the contribution from the charged particle background and other sources, a more realistic residual solar x-ray spectrum can be obtained by subtracting the background observed when the Sun is out of the FOV from the quiet Sun observations. However, this requires ignoring the periods when other bright sources contribute to the background. Sources that were bright enough to have a signal in XSM during the observing period were identified from the MAXI all-sky monitor data, and periods when such 13 sources were within the XSM FOV are ignored, resulting in an effective quiet Sun exposure of $2.15\times\mathrm{10^{6}}$ s. The resulting spectrum after subtracting the background is shown in the middle panel of Fig. 3(b) (Case 2), errors incorporate uncertainties on the observed quiet Sun spectrum and the background spectrum. We also consider Case 3, where the entire observed spectrum in Case 1 is assumed to arise from the background, and if the background were perfectly modeled, the resultant spectrum obtained is shown in the bottom panel of Fig. 3(b). This is obtained by empirically modeling the observed spectrum and subtracting the best-fit model from the observed spectrum.
@@
Data analysis. The expected number of photons per unit time per unit energy in a given energy bin $i$ of the XSM solely due to axion-photon conversion in the solar atmosphere is
$N^{i}_{\gamma,\,\rm exp}=\dfrac{R_{\odot}^{2}}{(\text{1 A.U.})^{2}}\,A_{\rm eff}(E_{i})\,\dfrac{d\phi_{a}}{dE_{a}}(E_{i})\,P_{a\gamma}(E_{i}),$ (8)
@@
where $E_{i}$ is the energy of the $i$-th energy bin, $A_{\rm eff}(E_{i})$ is the effective collecting area of the XSM in the $i$-th energy bin, and $R_{\odot}$ is the radius of the Sun. Given the observed number of counts per unit time per unit energy by the XSM, $N_{\gamma,\,obs}^{i}$, for a particular background subtraction scheme, we compute the likelihood assuming independent Gaussian statistic in each energy bin:
@@
$\mathcal{L}(g_{a\gamma\gamma};m_{a})\propto\prod_{i}\exp\left[-\tfrac{1}{2}\left(\frac{N_{\gamma,\exp}^{i}-N_{\gamma,\text{obs}}^{i}}{\sigma^{i}}\right)^{2}\right]$ (9)
@@
where $\sigma^{i}$ denotes the uncertainty in the observed counts in the $i$-th energy bin. To compute our limits we consider all the energy bins between $3.4$ keV and $15$ keV.
@@
To derive constraints in the $m_{a}-g_{a\gamma\gamma}$ plane, we fix the axion–electron coupling to $g_{aee}=1.3\times 10^{-13}$. For a given axion mass $m_{a}$, we adopt a uniform prior on the axion–photon coupling $g_{a\gamma\gamma}\in[10^{-14},\,10^{-8}]~\mathrm{GeV}^{-1}$. The posterior probability density $\mathcal{P}(g_{a\gamma\gamma}\,|\,\text{data})$ is obtained within a Bayesian framework using the PyMultiNest implementation of the MultiNest algorithm [Feroz:2008xx]. The 95% credible upper limit on $g_{a\gamma\gamma}$ is defined as the value enclosing 95% of the posterior probability. Repeating this procedure for axion masses in the range $m_{a}\in[10^{-12},\,10^{-3}]~\mathrm{eV}$ yields the corresponding mass-dependent upper limits on $g_{a\gamma\gamma}$.
In addition, we also derive constraints in the $g_{aee}-\,g_{a\gamma\gamma}$ plane by fixing the axion mass to $m_{a}=10^{-6}~\mathrm{eV}$. For $m_{a}\lesssim 10^{-4}~\mathrm{eV}$, the expected signal is effectively independent of $m_{a}$, so this choice is representative of the entire low-mass regime. The analysis is repeated while scanning over $g_{aee}$, yielding the corresponding 95% credible upper limits on $g_{a\gamma\gamma}$ as a function of $g_{aee}$.